# Fine-tuning foundational models to code diagnoses from veterinary health records

### v1

@@ 3.1 Input Formats
Preliminary experiments with the data input selection (Table 4) found that while the diagnosis and assessment sections of a visit record each individually provided some predictive capacity, the diagnosis portion was most useful. Note that the raw clinical text was necessarily truncated based on the maximum input length for each model (e.g., there was a 13% truncation rate when diagnosis and assessment were concatenated, in that order, for GatorTron). Diagnosis concatenated with assessment yielded marginally better results than either section alone did. Thus, diagnosis plus assessment were used as inputs for all models.
@@ 3.2 Performance Analysis
Figure 4: Code depth in Systematized Medical Nomenclature for Medicine - Clinical Terms (SNOMED-CT) hierarchy (relative to the most general “Clinical Finding” at level 0) plotted against GatorTron-based class F1. A larger value for depth indicated a higher degree of specificity (e.g., a leaf code such as “Poisoning Due to Rattlesnake Venom” (217659000) was at level 9). 530 codes could not be represented in this plot as they were older, inactive codes and thus were not incorporated into the newer version of the SNOMED-CT hierarchy used in this work. Exact match was not shown since it was calculated record-wise rather than class-wise. The same analysis was performed for other models and similar trends were observed. 95% confidence intervals for each level of depth are shown.
Performance was also observed to increase with larger amounts of fine-tuning data ($r_{F1}=0.85$, $r_{Precision}=0.72$, $r_{Recall}=0.90$, $r_{EM}=0.95$). All metrics followed a similar pattern with diminishing marginal returns on increased volumes of labeled fine-tuning data: using only 25% of all available labeled fine-tuning data yielded an 11.5% erosion of model performance, according to F1.
@@ 4 Discussion
The rapidly increasing availability of pre-trained LLMs [yang2022gatortron, wang2023optimized, han2023medalpaca, chen2023meditron70b, xie2024llama, hur2020domain, farrell2023petbert, bert2018, radford2019language, DBLP:journals/corr/abs-1907-11692] in recent years has greatly increased the ability of researchers to solve complex NLP tasks, including successful clinical coding with limited amounts of labeled fine-tuning data. The poor results (not reported here) of fine-tuning applied to several transformer models with randomly initialized weights highlighted the importance of significant pre-training for the task of clinical coding. GatorTron was pre-trained on the largest corpus of clinical data amongst all considered models (Table 1) and achieved the strongest performance after fine-tuning (Table 2). Additionally, performance on the downstream task (i.e., clinical coding) was not shown to be very sensitive to the domain of foundational LLM pre-training data. The tested human clinical LLMs (i.e., GatorTron, MedicalAI ClinicalBERT, and medAlpaca) and general domain LLMs (i.e., BERT, RoBERTa, and GPT-2) performed at least as well as veterinary clinical LLMs (i.e., VetBERT and PetBERT).
@@
Results from this study showed that the task of clinical coding was not very sensitive to model size. Relatively smaller models with respect to the total number of parameters (i.e., MedicalAI ClinicalBERT, VetBERT, PetBERT, BERT, RoBERTa, and GPT-2) performed comparably (6.2% worse F1 score, on average) to much larger models like GatorTron (Table 2). Further, performance was improved only relatively slightly by larger volumes of fine-tuning data, beyond a certain minimum. Accordingly, when the necessary computational resources to train relatively large LLMs are unavailable, smaller ones may be adequate for this task. A sensitivity analysis of the amount of fine-tuning data used for training demonstrated that similar results could be achieved using only a small fraction (e.g., 25% – 49,978 records) of the total amount of labeled data available (Figure 5). These two findings further reduce the barriers to application of automated clinical coding by demonstrating the potential sufficiency of relatively smaller LLMs and fine-tuning datasets. These are encouraging implications for situations in which limited computing resources, minimal labeled data, and reduced energy consumption are considerations [strubell2020energy].
@@
While a large volume of fine-tuning data, as was used in this study, may not always be necessary for a clinical coding task, some fine-tuning is clearly helpful. “Out-of-the-box” testing performed on all selected models generally yielded poor results (Table 3). The GPT-2 models were an exception as they performed much better than other models in the “out-of-the-box” setting. Nevertheless, even the GPT-2 models were improved by more expansive fine-tuning. These findings indicated that, building on pre-trained LLMs, the addition of fine-tuning with task-specific data is helpful in achieving improved results.
@@ 4.1 Performance Analysis
Errors made in diagnosis predictions can take many forms, and understanding the nature of these errors is important in recognizing the strengths and weaknesses of automated clinical coding and in informing future improvements. Understanding weaknesses can help prevent unsuitable application of developed tools. Several potential sources of error for the task of automated clinical coding include synonymous diagnoses recorded differently by clinicians (e.g., kidney disease and renal disease), non-synonymous diagnoses with similar names (e.g., gingival recession and gingivitis), and diagnoses that are inferred from histopathology or other reports that may not be recorded on a medical summary (e.g., endometritis is a histopathology diagnosis). Additionally, determination of the acceptable level of model performance for specific clinical tasks has not yet been established.
@@
No significant relationship between code specificity and predictability was observed (Figure 4). The prevalence of codes from many depths of the SNOMED-CT hierarchy indicates that clinicians and clinical coders at the CSU VTH choose to use diagnosis concepts with varying levels of specificity. Robust model performance across much of the hierarchy testifies to the flexibility of developed automatic coding tools with respect to the specificity of clinician input. This is important as practices for recording diagnoses can often vary on a clinician-by-clinician basis. The model’s generalizability across the SNOMED-CT hierarchy inspires confidence without more detailed diagnosis reporting standards being implemented.
A clear positive relationship between volume of data used for fine-tuning and model performance was observed (Figure 5). This multi-classification task is especially suited to large volumes of fine-tuning due to the large number of classes (7,739) involved and the relative rarity of a large number of the classes. However, model performance was found to be relatively more dependent when smaller amounts of fine-tuning data were used. Potentially useful model performance can be achieved with even a small fraction (e.g., 25%) of the CSU VTH fine-tuning dataset, with diminishing marginal returns for additional data. Certainly, the use of more labeled data for fine-tuning yields superior results, but having a labeled fine-tuning dataset smaller than the one from CSU VTH may still be useful. It should be noted that the sensitivity of this finding to factors like model size and domain of pre-training data has not yet been determined. This is a significant finding for sites without an existing large, well-curated, fine-tuning dataset where labeled data may have to be assembled from scratch.
@@ 4.3 Future Work
A major challenge to the growth of widespread use of AI in veterinary medicine is trust, both in clinical professionals and the public [omrani2022trust]. Evaluative tasks that help engender trust and confidence in model predictions will be included in future work in support of eventual applicability in a clinical setting. These tasks include increasing the explainability of the model (e.g., word saliency), demonstrating the generalizability of the model to applications beyond CSU (e.g., testing models against data from private practices or other academic veterinary institutions), performing in-depth error analyses by examining the source text associated with errors in collaboration with clinical experts (e.g., additional performance analyses), introducing new sources of textual data to model inputs (e.g., histopathology reports, patient history), and exploring more complex and potentially stronger [opitz1999popular] modeling processes (e.g., ensemble approaches).
The field of AI has been developing at an extraordinary pace. Since the inception of this project, several new foundational clinical language models have been released that could challenge the primacy of GatorTron. Yale University’s Me-LLaMA and the Swiss Federal Institute of Technology Lausanne’s MEDITRON have both been published since 2023. Due to computational limitations, neither of these models were directly considered for this project. However, future research in this field should leverage all possible foundational LLMs to remain cutting edge and determine the best solution for automating clinical diagnosis coding.
@@
Acknowledgements
@@
N. Saklou was supported by the National Institute of Health National Center for Advancing Translational Sciences (NCATS) U01TR002953-05. A. Kiehl, G. Strecker, and T. Webb were partially supported by NIH/NCATS Colorado CTSA Grant Number UM1 TR004399. Contents are the authors’ sole responsibility and do not necessarily represent official NIH views.
[…]

### v2

@@ 3.1 Input Formats
A sensitivity analysis with the data input selection for the best model (Table 4) found that while the diagnosis and assessment sections of a visit record each individually provided some predictive capacity, the diagnosis portion was most useful. Note that the raw clinical text was necessarily truncated based on the maximum input length for each model (e.g., there was a 16.7% truncation rate when diagnosis and assessment were concatenated, in that order, for GatorTron). Diagnosis concatenated with assessment yielded marginally better results than either section did alone.
@@ 3.2 Performance Analysis
Figure 4: Code depth in Systematized Medical Nomenclature for Medicine - Clinical Terms (SNOMED-CT) hierarchy (relative to the most general “Clinical Finding” at level 0) plotted against GatorTron-based class F1. A larger value for depth indicated a higher degree of specificity (e.g., a leaf code such as “Poisoning Due to Rattlesnake Venom” (217659000) was at level 9). 95% confidence intervals for each level of depth are shown.
Performance was also observed to increase with larger amounts of fine-tuning data ($r_{F1}=0.62$, $r_{Precision}=0.50$, $r_{Recall}=0.66$, $r_{EM}=0.74$; see Figure 5). All metrics followed a similar pattern with diminishing marginal returns on increased volumes of labeled fine-tuning data: using only 25% of all available labeled fine-tuning data yielded a 10.4% erosion of model performance, according to F1.
@@
Finally, model performance was aggregated by disease category based on the DAMNIT-V (Degenerative, Anomalous, Metabolic, Neoplastic, Nutritional, Inflammatory, Idiopathic, Toxic, Traumatic, Vascular) veterinary disease classification system. Model performance was largely robust across disease categories, with best performance observed for the metabolic disease category and worst performance observed for the nutritional disorder category.
Category N F1 Precision Recall
Other 174789 77.0 82.4 73.9
Neoplasm and/or hamartoma 64170 76.4 83.5 72.0
Inflammatory disorder 49261 82.9 88.0 79.7
Traumatic or non-traumatic injury 27710 70.2 79.3 65.2
Degenerative disorder 23132 81.4 88.6 76.2
Disorder of cardiovascular system 21946 78.9 86.7 74.4
Metabolic disease 11942 90.8 95.9 86.9
Congenital disease 7802 76.6 84.1 72.1
Poisoning 1044 79.7 84.9 75.6
Nutritional disorder 669 54.8 79.2 43.8
Table 5: GatorTron-based performance metrics aggregated by exclusive disease categories defined by the DAMNIT-V classification system for disease categories.
@@ 4 Discussion
The rapidly increasing availability of pre-trained LMs [120, 116, 75, 66, 119, 79, 71, 68, 103, 87, 76, 117] in recent years has greatly increased the ability of researchers to solve complex NLP tasks, including successful clinical coding with limited amounts of labeled fine-tuning data. The poor results (not reported here) of fine-tuning applied to several transformer models with randomly initialized weights highlighted the importance of significant pre-training for the task of clinical coding. GatorTron was pre-trained on the largest corpus of clinical data amongst all considered models (Table LABEL:table:Models) and achieved the strongest performance after fine-tuning (Table 2). Additionally, performance on the downstream task (i.e., clinical coding) was not shown to be very sensitive to the domain of foundational LM pre-training data. General pre-training data sources have been shown to contribute to robust performance across downstream tasks [88]. The tested human clinical LMs (GatorTron, MedicalAI ClinicalBERT, and medAlpaca) and general domain LMs (BERT, RoBERTa, GPT-2, and DeBERTa V3) performed at least as well as veterinary clinical LMs (VetBERT and PetBERT).
Several other models that deserve discussion were BERT Large, DeBERTa, and ModernBERT. BERT Large was fine-tuned under a slightly different training procedure that did not feature a dropout layer. This was due to convergence issues that arose only for this model. It is speculated that BERT Large may be undertrained [87] and may not be capable of fine-tuning on the sparse data involved in this project when dropout techniques are used. DeBERTa V3 and ModernBERT both claim to be more efficient alternative model architectures but were not appreciably faster than other models (0.21 and 0.41 hrs/epoch, respectively) and did not achieve stronger performance than other comparably-sized models.
@@
Results from this study showed that the task of clinical coding was not very sensitive to model size. Relatively smaller models with respect to the total number of parameters (MedicalAI ClinicalBERT, VetBERT, PetBERT, BERT, RoBERTa, GPT-2, DeBERTa V3, ModernBERT, Clinical ModernBERT) performed comparably (5.0% worse F1 score, on average) to much larger models like GatorTron (Table 2). Further, performance was improved only relatively slightly by larger volumes of fine-tuning data, beyond a certain minimum. Accordingly, when the necessary computational resources to train relatively large LMs are unavailable, smaller ones may be adequate for this task. A sensitivity analysis of the amount of fine-tuning data used for training demonstrated that similar results could be achieved using only a small fraction (e.g., 25% – 49,978 records) of the total amount of labeled data available (Figure 5). These two findings further reduce the barriers to application of automated clinical coding by demonstrating the potential sufficiency of relatively smaller LMs and fine-tuning datasets. These are encouraging implications for situations in which limited computing resources, minimal labeled data, and reduced energy consumption are considerations [106].
@@
While a large volume of fine-tuning data, as was used in this study, may not always be necessary for a clinical coding task, some fine-tuning is clearly helpful. “Out-of-the-box” testing performed on all selected models generally yielded poor results (Table 3). The GPT-2 models were an exception as they performed much better than other models in the “out-of-the-box” setting. Nevertheless, even the GPT-2 models were improved by more expansive fine-tuning. These findings indicated that, building on pre-trained LMs, the addition of fine-tuning with task-specific data is helpful in achieving improved results.
@@ 4.1 Performance Analysis
Errors made in diagnosis predictions can take many forms, and understanding the nature of these errors is important in recognizing the strengths and weaknesses of automated clinical coding and in informing future improvements. Understanding weaknesses can help prevent unsuitable application of developed tools. Several potential sources of error for the task of automated clinical coding include synonymous diagnoses recorded differently by clinicians (e.g., kidney disease and renal disease) and non-synonymous diagnoses with similar names (e.g., gingival recession and gingivitis). Additionally, determination of the acceptable level of model performance for specific clinical tasks has not yet been established.
@@
No significant relationship between code specificity and predictability was observed (Figure 4). The prevalence of codes from many depths of the SNOMED-CT hierarchy indicates that clinicians and clinical coders at the CSU VTH choose to use diagnosis concepts with varying levels of specificity. Robust model performance across much of the hierarchy testifies to the flexibility of developed automatic coding tools with respect to the specificity of clinician input. This is important as practices for recording diagnoses can often vary on a clinician-by-clinician basis. The model’s generalizability across the SNOMED-CT hierarchy inspires confidence without more detailed diagnosis reporting standards being implemented. 530 codes could not be represented in this analysis as they were older, inactive codes and thus were not incorporated into the newer version of the SNOMED-CT hierarchy used in this work. This analysis was performed for multiple models, and similar trends were observed.
A positive relationship between volume of data used for fine-tuning and model performance was observed (Figure 5). This multi-classification task is especially suited to large volumes of fine-tuning due to the large number of classes (7,739) involved and the relative rarity of a large number of the classes. However, model performance was found to be relatively more dependent when smaller amounts of fine-tuning data were used. Potentially useful model performance can be achieved with even a small fraction (e.g., 25%) of the CSU VTH fine-tuning dataset, with diminishing marginal returns for additional data. Certainly, the use of more labeled data for fine-tuning yields superior results, but having a labeled fine-tuning dataset smaller than the one from CSU VTH may still be useful. It should be noted that the sensitivity of this finding to factors like model size and domain of pre-training data has not yet been determined. This is a significant finding for sites without an existing large, well-curated, fine-tuning dataset where labeled data may have to be assembled from scratch.
Largely robust model performance was observed across high-level disease categories (Table 5. Understanding the strengths and weaknesses of developed models with respect to disease domains is crucial for responsible application. Only the nutritional disorder category demonstrated appreciably lower performance and was also the category with the fewest observations. While initial results showed encouraging robustness, many diagnosis codes used in this project have yet to be properly classified. The structure of the SNOMED-CT hierarchy does not allow for all codes of interest to be easily mapped to one of the selected high-level disease categories. For example, a manual mapping approach will be required to categorize diagnoses currently in the other category.
@@ 4.3 Future Work
A major challenge to the growth of widespread use of AI in veterinary medicine is trust, both in clinical professionals and the public [97]. Evaluative tasks that help engender trust and confidence in model predictions will be included in future work in support of eventual applicability in a clinical setting. An initial approach to model explainability used the integrated gradients method [107] to attribute predictive importance to input tokens. Similar methods have been implemented in other studies [92]. Future work will continue to expand on these efforts. Other tasks include (1) performing in-depth error analyses by examining the source text in collaboration with clinical experts (e.g., additional performance analyses), (2) demonstrating the generalizability of the model to applications beyond CSU (e.g., testing models against data from private practices or other academic veterinary institutions), (3) introducing new sources of textual data to model inputs (e.g., patient history, medication lists) and (4) selecting models with larger context windows [117, 62, 86], and exploring more complex and potentially stronger [98] modeling processes (e.g., ensemble approaches).
The field of AI has been developing at an extraordinary pace. Since the inception of this project, several new foundational clinical language models have been released that could challenge the primacy of GatorTron. Yale University’s Me-LLaMA [119] and the Swiss Federal Institute of Technology Lausanne’s MEDITRON [66] have both been published since 2023. Due to computational limitations, neither of these models were directly considered for this project. However, future research in this field should leverage all possible foundational LMs to remain cutting edge and determine the best solution for automating clinical diagnosis coding.
@@
Model Availability
The fine-tuned GatorTron model is freely available, subject to a DUA and a short CITI training, on PhysioNet.
Code Availability
All code used for data pre-processing and model training is freely available on GitHub.
Acknowledgments